40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1682     Entry time: Sat Jul 23 20:21:16 2016
Author: Antonio 
Type: DailyProgress 
Category: Mode matching 
Subject: South cavity Mode matching 

Summary
Following the previous entry the South path construction was at the point where mode matching was needed for the South Cavity.
We also needed to install a resonant photodiode for the PDH lock and start to prepare the optic in transmission.

Coonclusions

  • The light is resonating into the South Cavity with a visibility = (Vmax-Vmin)/Vmax = ~ 0.75. 
  • We also have the resonant PD for the light reflected from the cavity.
  • Camera and ISS PD in transmission of the South cavity have been aligned too.
  • All the cabeling are hooked, but they need to be tidied up.

Note:

  1. Lenses have been adjusted and slightly changed compared to the previous entry.
  2. The alignment of the cavity took me a while. The mounts for the steering mirrors on the
    periscope are very sensitive, here the alignment requires a bit of more attention. 
  3. All power supply and cables are connected. I just need to verify the HV power supply settings.

 

ELOG V3.1.3-